OpenStudy (anonymous):

simplify the trigonometric expression sec(theta) cos(theta) a.tan (theta) b.1 c.cot(theta) d.sin(theta)

OpenStudy (rajee_sam):

\[\sec \theta \cos \theta = \frac{ \cos \theta }{ \cos \theta } = 1\]

OpenStudy (anonymous):

thank you!